How much water should be added to 64 ounces of 10% salt solution to make a 2% salt solution?
Answer by stanbon(75887)

How much water should be added to 64 ounces of 10% salt solution to make a 2% salt solution?
---------------------
You start with 64 oz of solution and end up with (64+x)oz
----------------------
Equation:
salt = salt
0.10*64 = 0.02(64+x)
Multiply both sides by 100 to get:
10*64 = 2*64 + 2x
2x = 8*64
x = 8*32
x = 256 oz (amount of water needed to added to the 10% solution)
======================
Cheers,
Stan H.